# pgfplots.statistics

### v1

@@ Abstract
Index Terms:
Graph Neural Networks, Decentralized Control, Multi-Robot Systems, Robot Swarms
@@ I Introduction
Neural networks on graphs have been developed in the past decade for a variety of applications such as citation networks [1], classification of protein structures [9], and predicting power outages [10]. A fundamental difference between these applications and the collaboration of robot swarms is that the graph is generally static, whereas the robots are continuously moving, resulting in a dynamic and sparse communication graph. Moreover, control policies are executed on robots in real time with limited computational capabilities, which makes it imperative to provide an efficient framework for decentralized inference.
@@ II Navigation Control Problem
The navigation control problem, in the paper, considers a homogeneous team of $N$ mobile robots that needs to navigate through a $d$-dimensional environment $\mathcal{W}\subseteq\mathbb{R}^{d}$ to minimize the expected value of an application-defined cost function. We denote the set of robots by $\mathcal{V}=\{1,\ldots,N\}$, with the position of the $i$-th robot denoted by $\mathbf{p}_{i}(t)\in\mathcal{W}$ at time $t$. The state of the $i$-th robot at time $t$ is denoted by $\mathbf{x}_{i}(t)\in\mathbb{R}^{m}$, which in addition to robot positions may comprise velocities and other sensor measurements. The states of the multi-robot system and the control actions are collectively denoted by:
@@
The control problem can then be posed as an optimization problem, where the system incurs a cost given by a cost function $c(\mathbf{X}(t),\mathbf{U}(t))$ for state $X(t)$ and control action $U(t)$ taken at time $t$, and the goal is to find a policy $\Pi_{c}^{*}$ that minimizes the expected cost [4]:
@@ II-A Decentralized Navigation Control
In the decentralized navigation control problem, the robots take control actions based on their local information and the information communicated by their neighbors. Thus, we consider that the robots are equipped with a communication device for exchanging information with other robots that are within a given communication radius $r_{c}$. The problem can then be formulated on a communication graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the set of robots and $\mathcal{E}$ represents the communication topology of the robots. A robot $i$ can communicate with another robot $j$ if and only if their relative distance is less than a given communication radius $r_{c}$, i.e., an edge $e=(i,j)\in\mathcal{E}$ exists if and only if $\|\mathbf{p}_{i}-\mathbf{p}_{j}\|_{2}\leq r_{c}$. We assume bidirectional communication and, therefore, the communication graph is assumed to be undirected.
@@
Let $\frac{1}{\Delta t_{c}}$ be the frequency at which the robots communicate with each other, i.e., $\Delta t_{c}$ is the time required for a single communication to take place. Then the total information acquired by robot $i$ is given by:
@@ III Decentralized Graph Neural Networks
Graph Neural Networks (GNNs) [1] are layered information processing architecture that operate on a graph structure and make inferences by diffusing information through the graph. In the context of multi-robot systems, the graph is imposed by the communication graph, i.e., the graph nodes $\mathcal{V}$ represent the robots, and the edges represent the communication links $\mathcal{E}$, as discussed in Section II-A. In this paper, we focus on Graph Convolutional Neural Networks (GCNNs), a layered composition of convolution graph filters with point-wise nonlinearities ($\sigma$). The architecture is defined by $L$ layers, $K$ hops of message diffusion, and a shape shift operator $\mathbf{S}\in\mathbb{R}^{N\times N},\,N=\lvert\mathcal{V}\rvert$ that is a based on the communication graph. The elements $[\mathbf{S}]_{ij}$ can be non-zero only if $(i,j)\in\mathcal{E}$. The input to the GCNN is a collection of features $\mathbf{X}_{0}\in\mathbb{R}^{N\times d_{0}}$, where each element $\mathbf{x}_{i}$ is a feature vector for robot $i,\,\forall i\in\mathcal{V}$. The weight parameters for learning on GNNs are given by $\mathbf{H}_{lk}\in\mathbb{R}^{d_{(l-1)}\times d_{l}},\,\forall l\in\{1,\cdots,L\},\,\forall k\in\{1,\cdots,K\}$, where $d_{l}$ is the dimension of the hidden layer $l$ with $d_{0}$ as the dimension of the input feature vectors. The convolution graph filters are polynomials of the shape shift operator $\mathbf{S}$ with coefficients defined by the input and the weight parameters $\mathbf{H}_{lk}$. The output $\mathbf{Z}_{l}$ of the filter is processed by a point-wise non-linearity $\sigma$ to obtain the output of layer $l$, i.e., $\mathbf{X}_{l}=\sigma(\mathbf{Z}_{l})$. The final output of the GCNN is given by $\mathbf{X}_{L}$ and the entire network is denoted by $\Phi(\mathbf{X};\mathbf{S},\mathcal{H})$. Figure 3 shows an architecture with two layers.
@@
Here, $\mathcal{N}(i)$ is the set of neighbors of robot $i$, and $[\mathbf{S}]_{ij}$ is the $(i,j)$-th element of the shape shift operator $\mathbf{S}$. Since the value of $[\mathbf{S}]_{ij}$ is non-zero only if $(i,j)\in\mathcal{E}$, the computation of $(\mathbf{y}_{i})_{kl}$ only involves the features of the neighbors of robot $i$, i.e., the computation respects the locality of the communication graph. Thus, the robot $i$ needs to only receive $(\mathbf{y}_{j})_{(k-1)l}$ from its neighbors to compute $(\mathbf{y}_{i})_{kl}$, which makes the overall system decentralized. The collection of hidden feature vectors $(\mathbf{y}_{i})_{kl}$ forms the aggregated message [2] $\mathbf{Y}_{i}$ for robot $i$, which is precisely the information the robot $i$ needs to communicate to its neighboring robots.
@@
Note that the dimension of each vector in a row is same but the dimension across rows may be different, i.e., $\mathbf{Y}_{i}$ is a collection of matrices and not a proper tensor.
The overall algorithm for message aggregation and inference using a GNN is given in Algorithm 1. The output of the GNN is the output of the last layer $\mathbf{X}_{L}$. To completely diffuse a signal $\mathbf{x}_{i}$ across the network, each robot needs to exchange messages and execute Algorithm 1 for $KL$ times. This would require that the GNN message aggregation must be executed at a frequency that is $KL$ times higher than the frequency of perception. Depending on the application and the number of layers of the network, this may not be always feasible. However, due to the stability properties of the GNNs, they perform well even when the message aggregation is executed at a lower frequency.
@@ IV Asynchronous PAC Framework
Variable frequency: Asynchronous modules allow different subtasks to be executed at different frequencies. In particular, the GNN module can be executed at a higher frequency than the perception module, which has not been possible in prior work. Similarly, as is often the case, the communication and the low-level actuator controller modules can run at a higher frequency than the GNN message aggregation module. Generally, a GNN policy computes a high-level control action, which is then executed by a low-level controller for several time steps. Even in the case where the perception task is minimal, asynchronous execution allows the GNN module to perform several steps of message aggregation while the low-level controller is executing the previous control action.
@@
In prior work, the perception and GNN computations were considered to be executed synchronously, even if communication and low-level controller are asynchronous, affecting the computation time and the number of message aggregations that can be performed by the GNN module. This is mitigated in our proposed framework by allowing asynchronous execution. We now describe the different modules of the framework.
IV-A Perception
@@
The perception module is responsible for getting the sensor data and processing it to obtain the input features for the GNN. The module may also contain application-specific tasks such as semantic mapping, object detection, and localization, which makes the module computationally expensive. The entire perception module is typically executed at a low frequency in applications that require significant computation. In our specific implementation for the coverage control problem in Section V, we use a CNN to generate a low-dimensional feature vector for the GNN module.
IV-B Inter-Robot Communication
@@
IV-C GNN Message Aggregation and Inference
The GNN message aggregation module has two tasks: (1) generate messages to be communicated to neighboring robots in the next time step, and (2) perform inference for control actions. Our framework uses the aggregated GNN model, described in Section III. The system is fully decentralized, i.e., each robot has its own GNN model and the inference is performed locally. An important attribute of the aggregated GNN model is that the size of the messages is dependent on the number of layers in the network, and is independent of the number of robots. Thus, the system is highly scalable to large robot swarms. The module is executed at a higher frequency than the perception module.
@@ IV-D Low-Level Controller
The low-level controller interfaces the framework with the robot actuators. It receives the control action at the end of the computation of the GNN module and executes it for a pre-defined time interval. The controller is executed at a very high frequency to ensure that the control actions are reliably executed in real-time while correcting the control commands using a feedback loop. Since the implementation of the framework is designed to work with ROS2, any existing package for low-level control can be used with the framework.
V Evaluation on the Coverage Control Problem
(a) 1.25 Hz
(b) 1.67 Hz
(c) 2.5 Hz
(d) 5 Hz
Figure 4: Evaluation of coverage control algorithms, within our decentralized PAC framework, for four different operating frequencies (1.25 Hz to 5 Hz) of system evolution, which includes all perception tasks. The decentralized GNN, with CNN for perception, is trained using imitation learning with the ground truth generated using a clairvoyant Lloyd’s algorithm. The plots show the coverage cost, averaged over 30 environments, for 600 time steps and are normalized by the cost at the starting configuration of the system. Our approach significantly outperforms decentralized and centralized Lloyd’s algorithms.
(a) $\epsilon\sim\mathcal{N}(0,5^{2})$
(b) $\epsilon\sim\mathcal{N}(0,10^{2})$
(c) $\epsilon\sim\mathcal{N}(0,15^{2})$
(d) $\epsilon\sim\mathcal{N}(0,20^{2})$
Figure 5: Evaluation of coverage algorithms with a Gaussian noise $\epsilon$ added to the position of each robot, i.e., the sensed position $\mathbf{\bar{p}}_{i}=\mathbf{p}_{i}+\epsilon$. The perception module is executed at 2.5 Hz for all four cases. It is interesting that the performance of Lloyd’s algorithm increases slightly with noise, as this leads to discovering more of the IDF. However, the algorithms converge to a stable configuration after more number of steps. The decentralized GNN approach outperforms the two Lloyd’s algorithms in all cases, thereby demonstrating the robustness of our approach to noisy information.

### v2

# A Scalable Multi-Robot Framework for Decentralized and Asynchronous Perception-Action-Communication Loops

@@ I Introduction
Neural networks on graphs have been developed in the past decade for a variety of applications such as citation networks [1], classification of protein structures [9], and predicting power outages [10]. A fundamental difference between these applications and the collaboration of robot swarms is that the graph is generally static, whereas the robots are continuously moving, resulting in a dynamic and sparse communication graph. Moreover, control policies are executed on robots in real-time with limited computational capabilities, which makes it imperative to provide an efficient framework for decentralized inference.
@@ II Navigation Control Problem
The navigation control problem, in the paper, considers a homogeneous team of $N$ mobile robots that needs to navigate through a $d$-dimensional environment $\mathcal{W}\subseteq\mathbb{R}^{d}$ to minimize the expected value of an application-defined cost function. We denote the set of robots by $\mathcal{V}=\{1,\ldots,N\}$, with the position of the $i$-th robot denoted by $\mathbf{p}_{i}(t)\in\mathcal{W}$ at time $t$. The state of the $i$-th robot at time $t$ is denoted by $\mathbf{x}_{i}(t)\in\mathbb{R}^{m}$, which, in addition to robot positions, may comprise velocities and other sensor measurements. The states of the multi-robot system and the control actions are collectively denoted by:
@@
The control problem can then be posed as an optimization problem, where the system incurs a cost given by a cost function $c(\mathbf{X}(t),\mathbf{U}(t))$ for state $\mathbf{X}(t)$ and control action $\mathbf{U}(t)$ taken at time $t$, and the goal is to find a policy $\Pi_{c}^{*}$ that minimizes the expected cost [4]:
@@ II-A Decentralized Navigation Control
In the decentralized navigation control problem, the robots take control actions based on their local information and the information communicated by their neighbors. Thus, we consider that the robots are equipped with a communication device for exchanging information with other robots that are within a given communication radius $r_{c}$. The problem can then be formulated on a communication graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the set of robots and $\mathcal{E}$ represents the communication topology of the robots. A robot $i$ can communicate with another robot $j$ if and only if their relative distance is less than a given communication radius $r_{c}$, i.e., an edge $e=(i,j)\in\mathcal{E}$ exists if and only if $\|\mathbf{p}_{i}-\mathbf{p}_{j}\|_{2}\leq r_{c}$. We assume bidirectional communication, and therefore, the communication graph is assumed to be undirected.
@@
Let $\frac{1}{\Delta t_{c}}$ be the frequency at which the robots communicate with each other, i.e., $\Delta t_{c}$ is the time required for a single communication to take place. Then, the total information acquired by robot $i$ is given by:
@@ III Decentralized Graph Neural Networks
Graph Neural Networks (GNNs) [1] are layered information processing architecture that operate on a graph structure and make inferences by diffusing information through the graph. In the context of multi-robot systems, the graph is imposed by the communication graph, i.e., the graph nodes $\mathcal{V}$ represent the robots, and the edges represent the communication links $\mathcal{E}$, as discussed in Section II-A. In this paper, we focus on Graph Convolutional Neural Networks (GCNNs), a layered composition of convolution graph filters with point-wise nonlinearities ($\sigma$). The architecture is defined by $L$ layers, $K$ hops of message diffusion, and a graph shift operator $\mathbf{S}\in\mathbb{R}^{N\times N},\,N=\lvert\mathcal{V}\rvert$ that is a based on the communication graph. The elements $\mathbf{S}]_{ij}$ can be non-zero only if $(i,j)\in\mathcal{E}$. The input to the GCNN is a collection of features $\mathbf{X}_{0}\in\mathbb{R}^{N\times d_{0}}$, where each element $\mathbf{x}_{i}$ is a feature vector for robot $i,\,\forall i\in\mathcal{V}$. The weight parameters for learning on GNNs are given by $\mathbf{H}_{lk}\in\mathbb{R}^{d_{(l-1)}\times d_{l}},\,\forall l\in\{1,\cdots,L\},\,\forall k\in\{1,\cdots,K\}$, where $d_{l}$ is the dimension of the hidden layer $l$ with $d_{0}$ as the dimension of the input feature vectors. The convolution graph filters are polynomials of the graph shift operator $\mathbf{S}$ with coefficients defined by the input and the weight parameters $\mathbf{H}_{lk}$. The output $\mathbf{Z}_{l}$ of the filter is processed by a point-wise non-linearity $\sigma$ to obtain the output of layer $l$, i.e., $\mathbf{X}_{l}=\sigma(\mathbf{Z}_{l})$. The final output of the GCNN is given by $\mathbf{X}_{L}$, and the entire network is denoted by $\Phi(\mathbf{X};\mathbf{S},\mathcal{H})$. Figure 3 shows an architecture with two layers.
@@
Here, $\mathcal{N}(i)$ is the set of neighbors of robot $i$, and $\mathbf{S}]_{ij}$ is the $(i,j)$-th element of the graph shift operator $\mathbf{S}$. Since the value of $\mathbf{S}]_{ij}$ is non-zero only if $(i,j)\in\mathcal{E}$, the computation of $(\mathbf{y}_{i})_{kl}$ only involves the features of the neighbors of robot $i$, i.e., the computation respects the locality of the communication graph. Thus, the robot $i$ needs to only receive $(\mathbf{y}_{j})_{(k-1)l}$ from its neighbors to compute $(\mathbf{y}_{i})_{kl}$, which makes the overall system decentralized. The collection of hidden feature vectors $(\mathbf{y}_{i})_{kl}$ forms the aggregated message [2] $\mathbf{Y}_{i}$ for robot $i$, which is precisely the information the robot $i$ needs to communicate to its neighboring robots.
@@
Note that the dimension of each vector in a row is the same, but the dimension across rows may be different, i.e., $\mathbf{Y}_{i}$ is a collection of matrices and not a proper tensor.
The overall algorithm for message aggregation and inference using a GNN is given in Algorithm 1. The output of the GNN is the output of the last layer $\mathbf{X}_{L}$. To completely diffuse a signal $\mathbf{x}_{i}$ across the network, each robot needs to exchange messages and execute Algorithm 1 for $KL$ times. This would require that the GNN message aggregation must be executed at a frequency that is $KL$ times higher than the frequency of perception. Depending on the application and the number of layers of the network, this may not always be feasible. However, due to the stability properties of the GNNs, they perform well even when the message aggregation is executed at a lower frequency.
@@ IV Asynchronous PAC Framework
Variable frequency: Asynchronous modules allow different subtasks to be executed at different frequencies. In particular, the GNN module can be executed at a lower frequency than the perception module, which has not been possible in prior work. Similarly, as is often the case, the communication and the low-level actuator controller modules can run at a higher frequency than the GNN message aggregation module. Generally, a GNN policy computes a high-level control action, which is then executed by a low-level controller for several time steps. Even in the case where the perception task is minimal, asynchronous execution allows the GNN module to perform several steps of message aggregation while the low-level controller is executing the previous control action.
@@
In prior work, the perception and GNN computations were considered to be executed synchronously, even if communication and low-level controller are asynchronous, affecting the computation time and the number of message aggregations that can be performed by the GNN module. This is mitigated in our proposed framework by allowing asynchronous execution.
The system is physically realized across three domains: the robot (Section IV-B and Section V), the Ground Control Station (GCS) (Section IV-E) and the Communication network (Section VI). These three domains are staffed by an ensemble of ROS2 nodes run inside Docker containers deployed on the robots and the ground control station (GCS), while the Communication network serves as the incorporeal substrate. The use of containers is crucial to the software maintenance and the bring-up of a fleet of robots. We now describe the different modules and components of the framework.
IV-A Coordinate Frames
It is instructive to cast the system’s components in terms of their place in the set of four coordinate frames that are held in the system. These are: the mission frame $\mathcal{F}_{M}$, the North-East-Down (NED) local frame $\mathcal{F}_{L}$, the Front-Right-Down (FRD) body frame $\mathcal{F}_{L}$ and the global frame (Latitude and Longitude) $\mathcal{F}_{G}$. The complete set of frames is used on-board the vehicle and the ground control station deals only in coordinates referenced to $\mathcal{F}_{M}$ and $\mathcal{F}_{G}$. The key transform is between from $\mathcal{F}_{M}$ to $\mathcal{F}_{L}$, $\mathbf{T}_{M}^{L}$ and its inverse $\mathbf{T}_{L}^{M}=\textbf{inv}(\mathbf{T}_{M}^{L})$ because this is how positions and actions are transformed between LPAC to the flight controller and vice-versa; Figure 4.
Figure 4: Translation between $\mathcal{F}_{M}$ and $\mathcal{F}_{L}$ determined by the azimuth $\alpha$ and the distance $d$ between the mission origin $\mathcal{O}_{M}$ to the local origin $\mathcal{O}_{L}$. The phasor $d\angle\alpha$ is computed using the GPS coordinates of the vehicle at boot and the mission origin sent to it by the GCS.
IV-B LPAC
The LPAC node grants each robot autonomy and serves as the pilot; it operates exclusively in $\mathcal{F}_{M}$. It ingests sensory readings and positions as the robot traverses the IDF, as well as the perceptual embeddings $\mathbf{x}_{j}$ and positions $\prescript{M}{}{\mathbf{p}_{j}}$ in the mission frame received from neighboring robots ($j\in\mathcal{N}(i)$). From this input, LPAC produces the velocity $\prescript{M}{}{\mathbf{v}_{i}}$ that is delivered to the Offboard node to be transformed to $\prescript{L}{}{\mathbf{v}_{i}}$.
In our implementation, we set the Perception module to refresh the embedding $\mathbf{x}_{i}$ at a rate of $5\text{\,}\mathrm{Hz}$. We configure the GNN-based Communication module to query the message buffer manager, which holds positions and embeddings received from other robots $\mathbf{x}_{j},\prescript{M}{}{\mathbf{p}_{j}}]$, at a rate of $2.5\text{\,}\mathrm{Hz}$. Operating the Communications module at a lower frequency is a design decision that allows a greater window for aggregation. The robot’s velocity is read from the output buffer of the Action module and sent to the low-level flight controller at a rate of $10\text{\,}\mathrm{Hz}$.
IV-B1 Perception
@@
The perception module is responsible for getting the sensor data and processing it to obtain the input features for the GNN. The module may also contain application-specific tasks such as semantic mapping, object detection, and localization, which makes the module computationally expensive. The entire perception module is typically executed at a low frequency in applications that require significant computation. In our specific implementation for the coverage control problem in Section VII, we use a CNN to generate a low-dimensional feature vector for the GNN module.
IV-B2 Inter-Robot Communication
@@
IV-B3 GNN Message Aggregation and Inference
The GNN message aggregation module has two tasks: (1) generate messages to be communicated to neighboring robots in the next time step, and (2) perform inference for control actions. Our framework uses the aggregated GNN model, described in Section III. The system is fully decentralized, i.e., each robot has its own GNN model, and the inference is performed locally. An important attribute of the aggregated GNN model is that the size of the messages is dependent on the number of layers in the network, and is independent of the number of robots. Thus, the system is highly scalable to large robot swarms. The module is executed at a higher frequency than the perception module.
IV-C Offboard.
The offboard node manages the frame transformations between $\mathcal{F}_{M}$ and $\mathcal{F}_{L}$, and the state machine that, in turn, governs the mode of the robot. The frame transformations $\mathbf{T}_{M}^{L}$ and $\mathbf{T}_{L}^{M}$ require the computation of the phasor $d\angle\alpha$ with translation $d$ and azimuth $\alpha$, which can be found using a geodesic inverse function [15]. We can therefore define the translation for $\mathbf{T}_{L}^{M}$, $\mathbf{t}_{L}^{M}=[d_{x},d_{y},d_{z}]=d[\cos(\alpha),\sin(\alpha),0]$, where we assume that there is no difference in altitude between the frames. Using $\mathbf{t}_{L}^{M}$ and the heading $\theta$ which captures the difference in frame orientation between $\mathcal{F}_{L}$ and $\mathcal{F}_{B}$ we can express the transform in terms of its rotational and translational parts,
$\mathbf{T}_{L}^{M}=\left[\begin{array}[]{c@{}cc|c@{}}&\mathbf{R}_{z}\mathbf{R}_{x}&&\mathbf{t}_{L}^{M}\\ \hline\cr&0&&1\end{array}\right]\quad\text{where},$ (4)
$\mathbf{R}_{z}$ $=\begin{bmatrix}\cos\!\left(\tfrac{3\pi}{2}-\theta\right)&-\sin\!\left(\tfrac{3\pi}{2}-\theta\right)&0\\ \sin\!\left(\tfrac{3\pi}{2}-\theta\right)&\cos\!\left(\tfrac{3\pi}{2}-\theta\right)&0\\ 0&0&1\end{bmatrix}$, (5)
$\mathbf{R}_{x}$ $=\begin{bmatrix}1&0&0\\ 0&\cos\pi&-\sin\pi\\ 0&\sin\pi&\cos\pi\end{bmatrix}$
We compose the rotation $\mathbf{R}_{z}\mathbf{R}_{x}$ because the mission frame is established by pointing the robot’s body frame $x$-axis in the direction of the mission frame’s $y$-axis and record the heading $\theta$ (a convention chosen because it is easier to teleoperate a robot when it faces away from the operator). Therefore, we are first obliged to rotate a point in $\mathcal{F}_{L}$ about the $x$-axis by 180° followed by a rotation about the $z$-axis where the offset $3\pi/2$ comes from the fact that we originally defined $\theta$ to be the angle between $\mathcal{F}_{M}^{(y)}$ and $\mathcal{F}_{L}^{(x)}$.
The LPAC system is designed to maneuver vehicles on a 2D plane. Behavior along the vertical z-axis is completely independent of the LPAC control policy. Therefore, the Offboard node uses a simple proportional controller to drive each robot to a fixed altitude:
$v_{z}=-K_{p}(\mathbf{p}_{z}^{(d)}-\mathbf{p}_{c}^{(c)})$ (6)
Where $\mathbf{p}_{z}^{(d)}$ and $\mathbf{p}_{z}^{(c)}$ denote the desired position and current position along the vertical axis, respectively. We found that setting the proportional gain to $K_{p}=1$ worked well in practice.
The offboard node maintains the operational state-machine of the robot throughout an experiment. The machine receives messages from the GCS mission control node as the control input to progress the robot through a deployment. The offboard node controls all actuation while the robot is not in the Mission state. After takeoff altitude is reached, the offboard node begins transforming the velocities $\prescript{M}{}{\mathbf{v}_{i}}$ from the LPAC node to the low-level flight controller.
@@ IV-D Low-Level Controller
The low-level controller interfaces the framework with the robot actuators. It receives the control action at the end of the computation of the GNN module and executes it for a pre-defined time interval. The controller is executed at a very high frequency to ensure that the control actions are reliably executed in real time while correcting the control commands using a feedback loop. We use the PX4 flight controller which is an off-the-shelf autopilot software stack that abstracts the low-level actuation of the robot. It tracks both the local and global frames of the vehicle and therefore operates in $\mathcal{F}_{L}$ and $\mathcal{F}_{G}$. The PX4 interface is exposed to ROS2 through a MicroDDS pipeline [16] which enables reading the positions in the aformentioned frames and sending velocity commands to the robot. The origin $\mathcal{O}_{L}$ of $\mathcal{F}_{L}$ is determined at vehicle boot and the position in the frame is estimated using an Extended Kalman Filter (EKF) to fuse sensor readings. LPAC produces velocities $\prescript{M}{}{\mathbf{v}_{i}}$ in $\mathcal{F}_{M}$ which are transformed to the local frame, $\prescript{L}{}{\mathbf{v}_{i}}=\mathbf{T}_{M}^{L}\prescript{M}{}{\mathbf{v}_{i}}$. Likewise, positions from the flight controller $\prescript{L}{}{\mathbf{p}_{i}}$ can be transformed to the mission frame, $\prescript{M}{}{\mathbf{p}_{i}}=\mathbf{T}_{L}^{M}\prescript{L}{}{\mathbf{p}_{i}}$ to be used by LPAC. Since the implementation of the framework is designed to work with ROS2, any existing package for low-level control can be used with the framework.
IV-E Ground Control Station
IV-E1 Mission Control.
The Mission Control node is run on the GCS laptop and is the operator’s access panel to the state machine run in the offboard node on-board each robot. It includes safety features such as a geofence and the ability to severe autonomy and force the robots to hold position at the flip of a virtual switch. A routine experiment would involve the operator enabling the system, initiating takeoff, observing the mission and finally initiating a landing. The Mission Control node communicates with all robots on the network using a mission control ROS2 topic. The node also includes a field for specifying $\mathcal{O}_{M}$ in terms of the Latitude, Longitude and heading, which is then distributed to all robots on the network.
IV-E2 Mission Visualization.
A live view of the robots progress on coverage control is visualized by instantiating a local Coverage Control Simulator node on the GCS which receives positions from the robots. We remark that the simulator on the GCS does not influence the behaviour of the robots and is entirely for visualization purposes. The view of the agents overlaying the density function is rendered in RViz.
V Autonomous Aerial Vehicle Hardware
V-A Airframe and Compute
Figure 5: We use the ModalAI Starling 2 as our robot platform. The vehicle has onboard compute running both the flight controller and LPAC. Localization is accomplished using GNSS as well as inertial and barometric measurements. Inter-vehicle communication is handled by the onboard WiFi radios.
Autonomous aerial vehicles (AAV) serve as the main robotics platform for demonstrating the efficacy and robustness of our system. We use 10 of ModalAI’s Starling 2 quadrotor robots in each of our experiments. Our chosen platform is size weight and power (SWaP) limited; the $220\text{\,}\mathrm{mm}$ airframe weighs only $280\text{\,}\mathrm{g}$ [17]. From the factory, each Starling 2 is equipped with a Qualcomm QRB5165 processor mounted on a VOXL2 single board computer (SBC). The 64 bit ARM processor has 8 cores that are able to reach clock speeds of 2.84 GHz and is supported by 8GB of LPDDR5 memory [18]. LPAC inference and PX4 v1.14 autopilot run in tandem on the same on-board VOXL2 computer. The VOXL2 runs a modified version of Ubuntu 18.04 with the 4.19.125 kernel.
V-B Sensors
The focus of our experiments is on LPAC as a control policy and the interaction between agents via embeddings shared over the network. We operate under the assumption that the map of the environment has already been generated. Each robot must be able to localize itself in our mission frame $\mathcal{F}_{M}$ and maintain a fixed altitude. The simulated sensor coverage provided by each robot is constant and independent of altitude. Out of the box, the autopilot runs an extended Kalman filter (EKF) for sensor fusion. The vehicle is equipped with the uBlox-M10S GNSS receiver, QMC5883L compass, two TDK ICM-42688-P inertial measurement units (IMU) and two barometers: a Bosch BMP388 and a TDK ICP-10100. We configure the EKF to fuse GNSS, IMU and barometer measurements for localization. The Starling 2 includes several cameras, which we disable to save resources.
V-C Communication
LPAC relies on communication between agents. Each robot agent has an ALFA AWUS036EACS WiFi radio operating in the $5\text{\,}\mathrm{GHz}$ band. The radio has a typical power output of 17 dBm and has an integrated antenna. A single TP-Link AX1800 WiFi router manages traffic from vehicle to vehicle and between each vehicle and the ground control station (GCS). All radios in the system support the 802.11ac WiFi standard, enabling a maximum data-rate of 433 Mbps.
VI Network Configuration and Performance
The agents communicate over an 802.11-based wireless network operating in infrastructure mode with a shared access point (AP). To maximize coverage, the WiFi AP is centrally placed within the environment. For larger testing areas, configurations involving multiple routers connected via Ethernet have also been considered. This centralized setup simplifies coordination, as a single node manages all communication. However, such architectures can suffer from single points of failure and limited scalability as the number of agents grows. These trade-offs appear in any policy that involves inter-agent communication. Nonetheless, our architecture is agnostic to the structure of the communications network due to the inherent flexibility of the GNN. We use a centralized setup for the communications while computations remain decentralized. To support more scalable and fault-tolerant deployments, we have identified Doodle Labs radios [19] as a promising alternative for enabling fully distributed communication. Initial testing with these radios has yielded promising results, and they will be used in future implementations.
To better assess the quality of the network, we measured throughput using iPerf3 over a 10-second interval, varying the number of connected robots. The test configuration is shown in Figure 6, with the robots inside a dashed and continuous circle representing the iPerf3 server and client, respectively. Throughput was selected as the evaluation metric because it reflects overall network performance, capturing the effects of latency, jitter, and packet loss. The LPAC setup was disabled during these experiments to better evaluate the intrinsic network quality. For worst-case analysis, we focused on a drone located far from the router, and only considered one quadrant of the router’s coverage—assuming that only neighboring agents in that sector are relevant for the selected agent. As shown in Figure 7, network performance degrades as the number of connected robots increases. The bandwidth limitations of centralized network topologies highlight that a decentralized network is more amenable to teams with more than 10 agents.
Figure 6: Configuration for the throughput tests. An iPerf3 server was initialized in the drone with a dashed circle, while the drone with a continuous circle was configured as a client. Initially, only the first three robots in the diagonal were turned on. The other agents were turned on progressively.
Figure 7: Evolution of throughput measured in the network as the number of robots increases.
In initial experiments with 10 agents, we observed significant packet loss and latency, which hindered the smooth execution of the system. These issues highlighted the limitations of the default ROS2 middleware, DDS. Recent studies suggest that replacing DDS with the Zenoh protocol can improve both network performance and CPU utilization in wireless settings [20, 21]. To assess the potential benefits of this switch, we conducted a throughput test using the first three agents positioned along the diagonal in Figure 6. Each test ran for 60 seconds and was repeated 10 times under two configurations: using the default DDS and using Zenoh. To simulate network stress, the Docker containers on the robots were remotely restarted from the ground control station (GCS) during each interval. The average throughput across the 10 repetitions is reported in Table I. The results show that throughput is significantly higher when the Zenoh middleware is used. This enables more efficient data transmission by filtering and aggregating lower-level messages. Zenoh integrates seamlessly with ROS2, acting as a lightweight and high-performance alternative to the default DDS middleware. It enables communication across ROS2 nodes by bridging publish/subscribe and query/update paradigms in a unified protocol. This architecture reduces bandwidth usage and latency, making it especially suitable for dynamic and resource-constrained environments.
Uplink
throughput
Downlink
throughput
Without Zenoh $43.16\pm 7.56$ $42.66\pm 7.58$
With Zenoh $83.30\pm 1.45$ $82.89\pm 1.43$
TABLE I: Throughput achieved for a configuration with and without the Zenoh protocol.
VII Real-World Coverage Experiments